# Spacetime Spins: Statistical mechanics for error correction with stabilizer circuits

### v1

@@ A.1 Spacetime codes
The cumulant is an automorphism of $\mathcal{P}_{\text{st}}$ [delfosseSpacetime2023]; for $F,G\in\mathcal{P}_{\text{st}}$, it obeys $\overrightarrow{FG}=\overrightarrow{F}\overrightarrow{G}$. For a general $F\in\mathcal{P}_{\text{st}}$ that acts as $F_{\tau}\in\mathcal{P}_{\text{ph}}$ at each half-integer time $\tau$—that is, $F=\prod_{\tau=0.5}^{T-0.5}[F_{\tau}]_{,\tau}$—we thus have
@@
For a Pauli measurement $M_{\xi}\in\mathcal{P}_{\text{ph}}$ that occurs at $t_{\xi}$ in the circuit, let $o(M_{\xi})$ be its measurement outcome when the circuit is executed in the absence of errors. Depending on the circuit, this outcome may either be a deterministic value—that is, $M_{\xi}$ always gives the same outcome because the system is in its eigenstate when measured—or non-deterministic. If individual outcomes are non-deterministic, their product may still be deterministic if the product of the associated $M_{\xi}$ is an eigenoperator. For a set of measurement indices $\Xi_{k}=\{\xi\}$, if the product $\mathcal{O}_{k}=\prod_{\xi\in\Xi_{k}}o(M_{\xi})$ is deterministic, observing a different outcome indicates that an error occurred in the support of $\{M_{\xi}\}$. We use this property to define the stabilizers of the spacetime code,
@@
In a circuit that measures the syndromes of a QEC code, these $M_{(k)}$ take the form of two consecutive measurements of a stabilizer. These are equivalent to detector cells [mcewenRelaxing2023, derksDesigning2024, gottesmanOpportunities2022], which are generalizations of stabilizers to circuits: pairs of stabilizer measurements that signal an error whenever an outcome changes. The logical operators of the spacetime code are then the cumulants of the logical operators of the QEC code. Specifically, if a logical operator is $L(t)$ at timestep $t$ of a circuit (where the operator may evolve according to an automorphism in a Floquet code, for example), then the spacetime logical operator is the membrane
$\overrightarrow{L}=\prod_{t=0}^{T}[L(t)]_{,t+0.5}.$ (44)
If we have two logical operators $L_{1},L_{2}$ such that $\left\llbracket{L_{1}(t)},{L_{2}(t)}\right\rrbracket=-1$ for all $t$, then $\left\llbracket{\overrightarrow{L}_{1}},{\overrightarrow{L}_{2}}\right\rrbracket=-1$ iff $T$ is even [baconSparse2017]. [^15] If $T$ is odd, then the logical operator in Eq. (44) is in the gauge group and the code is trivial. If $T$ is odd, we may simply double it to achieve correct anticommutation. Since we are typically concerned with the $T\rightarrow\infty$ bulk behavior, we assume without loss of generality that $T$ is even.
These logical membranes are the bare logical operators of the spacetime subsystem code, in that they commute with all stabilizers and the gauge group, but are not in the gauge group. The dressed logical operators, however, typically take a simpler structure. In particular, $L(t_{1})]_{,t_{1}+0.5}$ and $L(t_{2})]_{,t_{2}+0.5}$ are gauge-equivalent, as the gauge operators encode propagations in spacetime. If $T$ is even, we thus have that $\overrightarrow{L}$ is gauge-equivalent to $L(t)]_{,t+0.5}$ for any $t\in[0,T]$. That is, the dressed spacetime logical operators include the original code’s logical operators on a single timeslice.
@@
When the Clifford circuit describes the syndrome measurement circuit of a stabilizer code, the two types of gauge generators discussed in the main text do indeed generate the spacetime gauge group. However, in a syndrome circuit of a subsystem code there are additional gauge freedoms not captured. Specifically, the gauge generators of the static subsystem code at each timestep need to be determined by considering the stabilizers measured by the circuit; these will add spacelike spacetime gauge operators [fuSubsystem2025]. One can check if all such operators have been added—or if they are necessary in the first place given an unknown Clifford circuit— by, for example, using Stim to search for low-weight undetectable errors [gidneyStim2021].
@@ A.2 SM mapping
The error-string model ultimately aims to relate the partition function of some quenched-disordered classical random-bond model to the probability of a class of logical errors. Specifically, define the logical coset
@@
For the total error string $\epsilon=\otimes_{i}\epsilon_{i}$, we therefore have
@@
Now, any errors $\epsilon,\epsilon^{\prime}\in E\mathcal{G}$ are related by a product of gauge operators, $\epsilon^{\prime}=g\epsilon$, $g\in\mathcal{G}$. If $\mathcal{G}=\braket{g_{1},g_{2},\ldots,g_{\gamma}}$ is generated by $\gamma$ gauge operators $g_{k}$ then we can represent $E\mathcal{G}$ by $\gamma$-bit binary vectors $\textsf{b}\in\textsf{B}_{\gamma}=\{0,1\}^{\gamma}$, as
$\epsilon_{\textsf{b}}=E\prod_{k=1}^{\gamma}g_{k}^{\textsf{b}_{k}}.$ (50)
@@
$\left\llbracket{\epsilon_{\textsf{b}}},{\alpha}\right\rrbracket$ $=\left\llbracket{E},{\alpha}\right\rrbracket\prod_{k=1}^{\gamma}\left\llbracket{g_{k}},{\alpha}\right\rrbracket^{\textsf{b}_{k}}$ (51)
@@
$\mathbb{P}(\overline{E})=\sum_{\textsf{b}\in\textsf{B}_{\gamma}}e^{-H_{E}(\textsf{b})},$ (52)
@@ A.2.1 Discrete Walsh transform
and we ignore the constant contribution. When $r=1$, $\cosh(P(\bm{\Sigma}))=\cosh|K_{1}\eta_{1}\sigma_{1}|$ is independent of $\eta_{1}\sigma=\pm 1$ and we are left with $H_{\text{eff}}=H_{0}$ up to constants. In this way, degree-$1$ spins and their sole interaction term can be removed from the Hamiltonian without affecting the partition function (up to an unimportant pre-factor).
@@
where $S$ is an index set that determines which components $\theta_{i}$ of $\bm{\theta}$ are included. A summation over all non-empty $S$ recovers $f(\bm{\Sigma})$ (up to constant terms) [berloffExact2025]
@@
When $r\geq 3$, the number of nonzero $F(S)$, and therefore the number of interactions in the effective Hamiltonian, is generally larger than $r$. We thus do not consider integrating out spins involved in more than two interactions; this also limits this procedure to independent $X$-$Z$ noise channels. Otherwise, contribution from $J(Y)$ terms increase the number of interactions each spin undergoes, and restrict this procedure to single-qubit gauge operators.
@@ B.1 Monte Carlo
The free-energy $F_{C_{\textsf{s}}L}$ is slow to compute due to the exponential function being dominated by high-energy but improbable configurations. We instead use the multi-step Bennett acceptance ratio [bennettEfficient1976, voterMonte1985]: to estimate $F_{C_{\textsf{s}}L}-F_{C_{\textsf{s}}}$, we decompose $L$ into $M$ “components”
$L_{\mu}=\prod_{m=1}^{\mu}Q_{m},\quad L_{0}=I,\quad L_{M}=L$ (76)
where each $Q_{m}\in\mathcal{P}_{\text{st}}$ is a (small) Pauli operator (e.g., a single-qubit Pauli). Then, we reduce the free-energy estimation to a product of smaller ratios
@@
We simulate $(M+1)$ $\beta$-ladders of $R$ replicas each with quenched disorder given by $C_{\textsf{s}},\,C_{\textsf{s}}L_{1},\,\ldots C_{\textsf{s}}L$ and compute
$\frac{\mathcal{Z}_{C_{\textsf{s}}L_{m}}}{\mathcal{Z}_{C_{\textsf{s}}L_{m-1}}}=\frac{\braket{f({C_{\textsf{s}}L_{m}};C_{\textsf{s}}L_{m-1}}_{\beta}}{\braket{f({C_{\textsf{s}}L_{m-1}};C_{\textsf{s}}L_{m}}_{\beta}},\quad f(E;E^{\prime})=\min\left\{1,e^{-\Delta\mathcal{E}(E;E^{\prime})}\right\}$ (78)
@@ B.2 Additional simulation details
Parallel tempering simulations were typically executed with 100020001000-2000 thermalization steps, $5000$ to $20000$ sweeps (over which the Bennett acceptance ratios were computed), and $10000$ to $20000$ quenched disorder realizations. 102010-20 temperature steps were used for the replicas. Population annealing typically used 2001000200-1000 thermalization steps, 100800100-800 population size, and $5000$ to $20000$ disorder realizations.
@@
Finally, intersection points of logical error rates were computed by taking sample data within the vicinity of the observed crossing point. Weighted least-squares regression (weighted by the sample errors) approximates each distance curve as a straight line, $y=m_{i}x+b_{i}$. The intersection point $(x_{c},y_{c})$ can be found by the overdetermined equation

### v2

@@ A.1 Spacetime codes
The cumulant is an automorphism of $\mathcal{P}_{\text{st}}$ [36]; for $F,G\in\mathcal{P}_{\text{st}}$, it obeys $\overrightarrow{FG}=\overrightarrow{F}\overrightarrow{G}$. For a general $F\in\mathcal{P}_{\text{st}}$ that acts as $F_{\tau}\in\mathcal{P}_{\text{ph}}$ at each half-integer time $\tau$—that is, $F=\prod_{\tau=0.5}^{T-0.5}[F_{\tau}]_{,\tau}$—we thus have
@@
For a Pauli measurement $M_{\xi}\in\mathcal{P}_{\text{ph}}$ that occurs at $t_{\xi}$ in the circuit, let $o(M_{\xi})$ be its measurement outcome when the circuit is executed in the absence of errors. Depending on the circuit, this outcome may either be a deterministic value—that is, $M_{\xi}$ always gives the same outcome because the system is in its eigenstate when measured—or non-deterministic. If an outcome is non-deterministic, the product of multiple outcomes may still be deterministic if the product of the associated $M_{\xi}$ is an eigenoperator. For a set of measurement indices $\Xi_{k}=\{\xi\}$, if the product $\mathcal{O}_{k}=\prod_{\xi\in\Xi_{k}}o(M_{\xi})$ is deterministic, observing a different outcome indicates that an error occurred in the support of $\{M_{\xi}\}$. We use this property to define the stabilizers of the spacetime code,
@@
In a circuit that measures the syndromes of a QEC code, these $M_{(k)}$ take the form of two consecutive measurements of a stabilizer. These are equivalent to detector cells [84, 38, 57], which are generalizations of stabilizers to circuits: pairs of stabilizer measurements that signal an error whenever an outcome changes. The logical operators of the spacetime code are then the cumulants of the logical operators of the QEC code. Specifically, if a logical operator is $L(t)$ at timestep $t$ of a circuit (where the operator may evolve according to an automorphism in a Floquet code, for example), then the spacetime logical operator is the membrane
$\overrightarrow{L}=\prod_{t=0}^{T-1}[L(t)]_{,t+0.5}.$ (44)
If we have two logical operators $L_{1},L_{2}$ such that $\left\llbracket{L_{1}(t)},{L_{2}(t)}\right\rrbracket=-1$ for all $t$, then $\left\llbracket{\overrightarrow{L}_{1}},{\overrightarrow{L}_{2}}\right\rrbracket=-1$ iff $T$ is odd [7]. If $T$ is even, then the logical operator in Eq. (44) is in the gauge group and the code is trivial. In such a case, we may simply repeat the circuit to achieve correct anticommutation.
These logical membranes are the bare logical operators of the spacetime subsystem code, in that they commute with all stabilizers and the gauge group, but are not in the gauge group. The dressed logical operators typically take a simpler structure. In particular, $L(t_{1})]_{,t_{1}+0.5}$ and $L(t_{2})]_{,t_{2}+0.5}$ are gauge-equivalent, as the gauge operators encode propagations in spacetime. If $T$ is odd, we thus have that $\overrightarrow{L}$ is gauge-equivalent to $L(t)]_{,t+0.5}$ for any $t\in[0,T-1]$. That is, the dressed spacetime logical operators include the original code’s logical operators on a single timeslice.
@@
When the Clifford circuit describes the syndrome measurement circuit of a stabilizer code, the two types of gauge generators discussed in the main text do indeed generate the spacetime gauge group. However, in a syndrome circuit of a subsystem code there are additional gauge freedoms not captured. Specifically, the gauge generators of the static subsystem code at each timestep need to be determined by considering the stabilizers measured by the circuit; these will add spacelike spacetime gauge operators [50]. One can check if all such operators have been added—or if they are necessary in the first place given an unknown Clifford circuit— by, for example, using Stim to search for low-weight undetectable errors [54].
@@ A.2 SM mapping
The SM model ultimately aims to relate the partition function of some quenched-disordered classical random-bond model to the probability of a class of logical errors. Specifically, define the logical coset
@@
For a generic multi-qubit error $\epsilon=\otimes_{i}\epsilon_{i}$, we therefore have
@@
Now, any errors $\epsilon,\epsilon^{\prime}\in E\mathcal{G}$ are related by a product of gauge operators, $\epsilon^{\prime}=g\epsilon$, $g\in\mathcal{G}$. If $\mathcal{G}=\braket{g_{1},g_{2},\ldots,g_{m}}$ is generated by $m$ gauge operators $g_{k}$ then we can represent $E\mathcal{G}$ by $m$-bit binary vectors $\textsf{b}\in\textsf{B}_{m}=\{0,1\}^{m}$, as
$\epsilon_{\textsf{b}}=E\prod_{k=1}^{m}g_{k}^{\textsf{b}_{k}}.$ (50)
@@
$\left\llbracket{\epsilon_{\textsf{b}}},{\alpha}\right\rrbracket$ $=\left\llbracket{E},{\alpha}\right\rrbracket\prod_{k=1}^{m}\left\llbracket{g_{k}},{\alpha}\right\rrbracket^{\textsf{b}_{k}}$ (51)
@@
$\mathbb{P}(\overline{E})=\sum_{\textsf{b}\in\textsf{B}_{m}}e^{-H_{E}(\textsf{b})},$ (52)
@@ A.2.1 Discrete Walsh transform
and we ignore the constant contribution. When $r=1$, $\cosh[P(\bm{\Sigma})]=\cosh|K_{1}\eta_{1}\sigma_{1}|$ is independent of $\eta_{1}\sigma=\pm 1$ and we are left with $H_{\text{eff}}=H_{0}$ up to constants. In this way, degree-$1$ spins and their sole interaction term can be removed from the Hamiltonian without affecting the partition function (up to an unimportant pre-factor).
@@
where $S$ is an index set that determines which components $\theta_{i}$ of $\bm{\theta}$ are included. A summation over all non-empty $S$ recovers $f(\bm{\Sigma})$ (up to constant terms) [13]
@@
When $r\geq 3$, the number of nonzero $F(S)$, and therefore the number of interactions in the effective Hamiltonian, is generally larger than $r$. We thus do not consider integrating out spins involved in more than two interactions; this also limits this procedure to independent $X$-$Z$ noise channels. Otherwise, contribution from $K(Y)$ terms increase the number of interactions each spin undergoes, and restrict this procedure to single-qubit gauge operators.
A.3 Additional spin diagram simplifications
In addition to the two simplification rules discussed in the main text, cf. Fig 3, a third rule can also be described that involves changing the basis of gauge generators. That is, for two gauge generators $g_{1},g_{2}$, the mapping $g_{2}\mapsto g_{2}^{\prime}=g_{1}g_{2}$ preserves $\mathcal{G}$. Diagrammatically, we draw this by updating the interactions associated with the spin for $g_{2}$:
This rule may be useful for re-arranging bonds to reveal simpler or more informative spin diagrams. For example, using this rule we can investigate the differences between a SWAP gate implemented natively (or equivalently, as a phenomenological representation) as in Fig. 4, and a SWAP gate implemented via three CNOT gates:
(The simplifying rules used in each step are shown in parentheses.) Unlike the native implementation, this spin diagram includes an additional spin and interaction that highlights the potential for hook errors to spread between the two wires. Spin diagrams, in conjunction with the three simplification rules, thus serve as a useful tool for understanding the propagation of errors and the differences between implementations of circuit operations.
A.4 Additional figures
Figure A.1: MWPM logical error rates for two repetition codes with standard syndrome extraction circuits. (Left) Subject to one transversal CNOT, with N=4d2N=4d-2, T=4d1T=4d-1. Unlike Fig. 12 with T=16d1T=16d-1, no threshold is observed at this reduced circuit depth. (Right) Subject to $d$ CNOTs, with N=4d2N=4d-2 and T=16d1T=16d-1. Unlike the ML decoder in Fig. 13, the MWPM decoder has no threshold.
@@ B.1 Monte Carlo
The free energy $F_{C_{\textsf{s}}L}$ is slow to compute due to the exponential function being dominated by high-energy but improbable configurations. We instead use the multi-step Bennett acceptance ratio [12, 117]: to estimate $F_{C_{\textsf{s}}L}-F_{C_{\textsf{s}}}$, we decompose $L$ into $M$ “components” $Q_{1},Q_{2},\ldots,Q_{M}$, where each $Q_{\mu}\in\mathcal{P}_{\text{st}}$ is a (small) Pauli operator (e.g., a single-qubit Pauli). They satisfy
$L_{m}=\prod_{\mu=1}^{m}Q_{\mu},\quad L_{0}=I,\quad L_{M}=L.$ (76)
Then, we reduce the free-energy estimation to a product of smaller ratios
@@
We simulate $(M+1)$ $\beta$-ladders of $R$ replicas each with quenched disorder given by $C_{\textsf{s}},\,C_{\textsf{s}}L_{1},\,\ldots,C_{\textsf{s}}L$ and compute
$\frac{\mathcal{Z}_{C_{\textsf{s}}L_{m}}}{\mathcal{Z}_{C_{\textsf{s}}L_{m-1}}}=\frac{\braket{f({C_{\textsf{s}}L_{m}};C_{\textsf{s}}L_{m-1})}_{\beta}}{\braket{f({C_{\textsf{s}}L_{m-1}};C_{\textsf{s}}L_{m})}_{\beta}},\quad f(E;E^{\prime})=\min\left\{1,e^{-\Delta\mathcal{E}(E;E^{\prime})}\right\}$ (78)
@@ B.2 Additional simulation details
Parallel tempering simulations were typically executed with 100020001000-2000 thermalization steps, $5000$ to 20 00020\,000 sweeps (over which the Bennett acceptance ratios were computed), and 10 00010\,000 to 50 00050\,000 quenched disorder realizations. 102010-20 temperature steps were used for the replicas. Population annealing typically used 2001000200-1000 thermalization steps, 100800100-800 population size, and $5000$ to 20 00020\,000 disorder realizations.
@@
Finally, intersection points of logical error rates (for thresholds) were computed by taking sample data within the vicinity of the observed crossing point. Weighted least-squares regression (weighted by the sample errors) approximates each distance curve as a straight line, $y=m_{i}x+b_{i}$. The intersection point $(x_{c},y_{c})$ can be found by the overdetermined equation